# DEM: A Method for Certifying Deep Neural Network Classifier Outputs in Aerospace

### v1

@@ 2 Related Work
A recent work [14] proposed an approach for assessing DNN robustness using an alternative Functional Hazard Analysis (FHA) method. This method is similar to the one required by ARP-4754, but does not quite address the requirements in the guidelines.
Another recent study [11] developed a comprehensive framework of principles for DNN certification, which is aligned with ARP-4754 — but which does not include a concrete method or tool that can be used.
While ARP-4754 focuses on system level certification, the influential DO-178 guidelines for airborne systems focus on software certification [10]. These guidelines define 5 levels (A-E) for software failure conditions. Initial work has discussed certifying DNNs to level D [5, 6] or C [4, 7, 4], but to date none have reached Level A — the most severe level, required where failures may lead to fatalities, which is the case in aerospace. We hope that our approach can be used to satisfy Level A requirements in the future. DNN Robustness. An alternative approach to the one proposed here, which has received significant attention, is the formal certification of DNN robustness. The idea is to conclude, a-priori, that a DNN is safe to use, at least for certain regions of its input space.
@@ 5.1 Evaluating the Recall-Oriented Calibration Algorithm
Class
Genuine Recall
Adversarial Recall
#Samples RES|VGG
DEM LID #Samples RES|VGG DEM LID
Airplane
180|177
63%|76% 81%|88% 393|125 92%|86% 54%|30%
Automotive
190|185
65%|65% 80%|91% 290|80 94%|87% 43%|52%
Bird
169|160
79%|67% 80%|84% 441|159 91%|91% 44%|35%
Cat
149|146
51%|50% 74%|85% 547|246 91%|89% 52%|31%
Deer
179|176
80%|79% 76%|89% 526|186 93%|97% 47%|37%
Dog
164|135
56%|60% 74%|87% 465|183 91%|92% 44%|30%
Frog
183|174
99%|91% 76%|84% 452|138 99%|97% 54%|34%
Horse
183|174
65%|74% 84%|88% 404|125 88%|87% 44%|38%
Ship
188|185
88%|91% 79%|92% 260|64 96%|92% 57%|56%
Truck
185|184
86%|80% 83%|87% 331|100 95%|95% 45%|44%
@@ 5.2 Evaluating the Precision-Oriented Calibration Algorithm
Class
Genuine Examples
Adversarial Inputs
#Samples RES|VGG
Precision Recall #Samples RES|VGG Precision Recall
Airplane
180|177
88%|85% 61%|80% 393|125 80%|81% 68%|86%
Automotive
190|185
91%|82% 68%|73% 290|80 77%|76% 91%|84%
Bird
169|160
89%|87% 79%|78% 441|159 81%|81% 91%|79%
Cat
149|146
91%|83% 41%|53% 547|246 77%|73% 69%|55%
Deer
179|176
90%|85% 82%|91% 526|186 84%|90% 91%|84%
Dog
164|135
88%|80% 50%|68% 465|183 74%|75% 74%|53%
Frog
183|174
88%|83% 100%|99% 452|138 100%|99% 86%|80%
Horse
183|174
85%|84% 64%|79% 404|125 75%|80% 80%|85%
Ship
188|185
90%|84% 95%|94% 260|64 94%|93% 90%|83%
Truck
185|184
89%|87% 91%|91% 331|100 91%|91% 88%|86%
@@ 6 Discussion and Future Work
Notably, in certain models and specific categories, DEMachieves flawless results of 100%, suggesting its potential suitability for adoption in the aerospace industry.

### v2

@@ 2 Related Work
A recent work [14] proposed an approach for assessing DNN robustness using an alternative Functional Hazard Analysis (FHA) method. This method is similar to the one required by ARP-4754, but does not quite address the requirements in the guidelines. Another recent study [11] developed a comprehensive framework of principles for DNN certification, which is aligned with ARP-4754 — but which does not include a concrete method or tool that can be used.
While ARP-4754 focuses on system level certification, the influential DO-178 guidelines for airborne systems focus on software certification [10]. These guidelines define 5 levels (A-E) for software failure conditions. Initial work has discussed certifying DNNs to level D [5, 6] or C [4, 7, 4], but to date none have reached Level A — the most severe level, required where failures may lead to fatalities, which is the case in aerospace. We hope that our approach can be used to satisfy Level A requirements in the future.
DNN Robustness. An alternative approach to the one proposed here, which has received significant attention, is the formal certification of DNN robustness. The idea is to conclude, a-priori, that a DNN is safe to use, at least for certain regions of its input space.
@@ 4.2 Dataset Preparation and Calibration
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 5.1 Evaluating the Recall-Oriented Calibration Algorithm
{tblr}
[…]
#Samples RES|VGG DEM LID #Samples RES|VGG DEM LID
Airplane 180|177 63%|76% 81%|88% 393|125 92%|86% 54%|30%
Automotive 190|185 65%|65% 80%|91% 290|80 94%|87% 43%|52%
Bird 169|160 79%|67% 80%|84% 441|159 91%|91% 44%|35%
Cat 149|146 51%|50% 74%|85% 547|246 91%|89% 52%|31%
Deer 179|176 80%|79% 76%|89% 526|186 93%|97% 47%|37%
Dog 164|135 56%|60% 74%|87% 465|183 91%|92% 44%|30%
Frog 183|174 99%|91% 76%|84% 452|138 99%|97% 54%|34%
Horse 183|174 65%|74% 84%|88% 404|125 88%|87% 44%|38%
Ship 188|185 88%|91% 79%|92% 260|64 96%|92% 57%|56%
Truck 185|184 86%|80% 83%|87% 331|100 95%|95% 45%|44%
@@ 5.2 Evaluating the Precision-Oriented Calibration Algorithm
{tblr}
[…]
#Samples RES|VGG Precision Recall #Samples RES|VGG Precision Recall
Airplane 180|177 88%|85% 61%|80% 393|125 80%|81% 68%|86%
Automotive 190|185 91%|82% 68%|73% 290|80 77%|76% 91%|84%
Bird 169|160 89%|87% 79%|78% 441|159 81%|81% 91%|79%
Cat 149|146 91%|83% 41%|53% 547|246 77%|73% 69%|55%
Deer 179|176 90%|85% 82%|91% 526|186 84%|90% 91%|84%
Dog 164|135 88%|80% 50%|68% 465|183 74%|75% 74%|53%
Frog 183|174 88%|83% 100%|99% 452|138 100%|99% 86%|80%
Horse 183|174 85%|84% 64%|79% 404|125 75%|80% 80%|85%
Ship 188|185 90%|84% 95%|94% 260|64 94%|93% 90%|83%
Truck 185|184 89%|87% 91%|91% 331|100 91%|91% 88%|86%